ünif uniform random variable 1) Draw the graphs of the p.d.f. of the following distributions (a) The standard normal p.d.f (b) The normal pdf with ? = 50, ? = 10 (c) The uniform p.d.f. over interval [10, 20] (d) The exponential P.d.f with parameter ? 4. 2) Illustrating the central limit theorem. Let X be a random variable having the uniform distribution over the interval [6, 12] Denote by X1, X. X3, a sequence of independent random variables with the same distribution as X. Define the sample mean by X The central limit theorem applied to this particular case implies that the probability distribution of 7 converges to the standard normal distribution for certain values of ? and ? (a) For what values of and ? does the convergence hold? (This is to be done by hand.) (b) For each of the four values n-2, 5, 10, and 20 do the following: Obtain 10000 independent values of the sample mean random variable X, where the X are generated from the uniform distribution over the interval [6, 12]; then create a histogram of for eachn. Discuss what you notice for each case on n.I must use R Program to solve them. Please help! Thank you
Answer #1

1)

a) The graph of N\left ( 0,1 \right ) is drawn below using R.

R code given below.

x <- seq(-5, 5, length=1000)
y <- dnorm(x, mean=0, sd=1)
plot(x, y, type="l", lwd=1)

b)The graph of N\left ( 50,10^2 \right ) is drawn below using R.


R code given below.0.03 0.04 0.01 0.02 0.00 3 8

x <- seq(10, 90, length=1000)
y <- dnorm(x, mean=50, sd=10)
plot(x, y, type="l", lwd=1)

c) The graph of Unif\left ( 10,20 \right ) is drawn below using R.

0.00 0.02 0.04 0.06 0.08 0.10 8

R code given below.

x <- seq(5, 25, length=1000)
y <- dunif(x, 10, 20)
plot(x, y, type="l", lwd=1)

d) The graph of Exp\left ( \lambda =4\right ) is drawn below using R.

0 9 8 7 6 5 60-ag 60-ag 60-at 60-az oo+90

R code given below.

x <- seq(5, 10, length=1000)
y <- dexp(x, 4)
plot(x, y, type="l", lwd=1)